Naina runs a business that only sells hamburgers in a building she owns. Every month, they sell 6,000 hamburgers at $5 per hamburger. She spends $2 per hamburger on supplies (meat, lettuce, bread, etc.). She also pays John and Rajeev each $2500 per month to work at the restaurant. Finally, utilities cost $500 per month. Susana works full-time at the restaurant and keeps the accounting profits for herself. What is the accounting profit of the business? Susana could rent out her building for $5000 per month. She can also make $6000 per month as an accountant. Based on this, what is the economic profit of her business? 


1
Expert's answer
2021-05-04T12:15:01-0400

the accounting profit


"\pi=TR-TC"

"\pi=6000\times5-(2\times2500+500+2\times6000)=12500"

the economic profit


"12500-11000=1500"